### v2

# Why Study Spherical Convexity of Non-Homogeneous Quadratic Functions? What Makes It Surprising?

Abstract
This paper establishes necessary, sufficient, and equivalent conditions for the spherical convexity of non-homogeneous quadratic functions. By examining criteria for determining spherical convexity, we identified unique properties that differentiate spherically convex quadratic functions from their geodesically convex counterparts in both hyperbolic and Euclidean spaces. Since spherically convex functions over the entire sphere are constant, our analysis focuses on proper spherically convex subsets of the sphere. Our primary results concern non-homogeneous quadratic functions on the spherically convex set of unit vectors with positive coordinates. We also extend our findings to more general spherically convex sets. Additionally, the paper explores special cases of non-homogeneous quadratic functions where the defining matrix is of a specific type, such as positive, diagonal, or a Z-matrix. This study not only provides useful criteria for spherical convexity but also reveals surprising characteristics of spherically convex quadratic functions, contributing to a deeper understanding of convexity in spherical geometries.
1 Introduction
This study was prompted by a reviewer of the paper [14], who suggested analyzing the geodesic convexity of non-homogeneous quadratic functions on the sphere as a natural counterpart to the same problem on a hyperbolic manifold.
Initially, we anticipated implementing similar methodologies to those used in the aforementioned paper. However, we found that the approaches employed in our study needed to diverge significantly from those applied to the analogous problem on the hyperbolic manifold.
This disparity arises from our examination of the geodesic convexity of non-homogeneous convex functions across the entire hyperbolic manifold, a problem that becomes trivial on the sphere. Indeed, such a task presents no difficulty on the sphere due to all its geodesics being closed, rendering geodesically convex functions on the sphere constant. Consequently, we had to confine our investigation to geodesically convex subsets of the sphere, which we termed ”spherically convex”. As a result, the problem became significantly more challenging since linear coordinate transformations, akin to those employed in [14], failed to leave the considered spherically convex set invariant. The most intuitive spherically convex set is the collection of points on the sphere with nonnegative coordinates, i.e., the intersection of the sphere with the nonnegative orthant. While much of our study has focused on the interior of this set, we have also derived results for a general spherically convex set, defined as the intersection of the sphere with a pointed convex cone.
Motivations for studying the geodesic convexity of non-homogeneous spherically convex quadratic functions are outlined below:
  1. 1.
    Generally, the utility of exploring geodesic convexity mirrors that of Euclidean convexity.
    1. (a)
      A local minimizer of a geodesic convex function is also a global minimizer, and strict geodesic convexity ensures a unique minimizer.
    2. (b)
      In general optimization algorithms on Riemannian manifolds exhibit better convergence properties and superior performance when optimizing geodesically convex functions.
  2. 2.
    Numerous research papers and books (e.g., [24, 8, 11, 12, 29, 17, 32, 37, 38, 2, 7, 23, 39, 3, 35, 1]) have leveraged the concept of geodesic convexity of functions on manifolds to establish various optimization results and analyze algorithmic convergence.
    It is imperative to exercise caution when employing theories based on geodesic convexity without a comprehensive understanding of its nuances.
    For instance, several publications in reputable journals have made incorrect assumptions regarding geodesic convexity on Hadamard manifolds, erroneously extending conditions that hold only under zero sectional curvature, as detailed in [22]. Determining the geodesic convexity of a function on a manifold often presents formidable challenges, even for seemingly straightforward cases involving simple functions on manifolds of constant nonzero curvature, such as quadratic functions.
    The seemingly basic task of discerning whether a quadratic function on a manifold of constant curvature is geodesically convex highlights the complexity inherent in this endeavor. With the burgeoning interest in manifold optimization, it is paramount to establish a robust foundation. If the research community in manifold optimization struggles to address fundamental inquiries, such as the geodesic convexity of elementary functions on basic manifolds, it raises significant concerns about the field’s credibility.
  3. 3.
    Several important problems and applications are related to minimizing non-homogeneous quadratic functions on the sphere (see [34, 19] together with the corresponding references [4, 16, 18, 26, 27, 33, 5, 6, 20, 25, 31, 36] listed in their introduction). These problems are related to trust region algorithms, regularization methods for ill-posed problems, and seismic inversion problems, as detailed in [34, 19]. The convergence and performance of possible algorithms addressing these problems can profit from the geodesic convexity displayed by such functions.
While our study may appear peripheral in the domain of manifold optimization, the aforementioned motivations underscore the indispensable significance of this topic.
For clarity, we refer to the intersection of the positive orthant with the sphere as the ”spherical positive orthant.” The key findings of the paper are summarized as follows:
  1. 1.
    An affine function on a spherically convex set, defined by a proper cone, is spherically convex if and only if its linear part is defined by a vector in the cone’s polar.
  2. 2.
    While our previous work (Ferreira & Nemeth, 2019) demonstrated that all convex, homogeneous quadratic functions on the spherical positive orthant are constant, the class of non-homogeneous quadratic functions on this domain is surprisingly vast.
  3. 3.
    Contrary to the case in hyperbolic space, where any non-homogeneous geodesically convex function has a geodesically convex homogeneous quadratic part, this property doesn’t extend to the spherical nonnegative orthant. Moreover, for any symmetric matrix $A\in\mathbb{R}^{n\times n}$, there exist infinitely many non-homogeneous spherically convex quadratic functions with $A$ as the matrix of their homogeneous quadratic part. The spherical convexity of these functions depends on the size of the components of the vector $b$ defining their first-order homogeneous component. The largest component of $b$ is unique if the domain of the spherically convex function is the positive orthant, being the component with the smallest modulus, while the other components of $b$ are negative.
  4. 4.
    The function described in item 3 can exhibit spherically convex behavior even with a positive largest component of $b$, as demonstrated in our work. Determining the smallest upper bounds of the components of $b$ for spherically convex behavior poses a challenging question, possibly involving unknown relationships between these components as part of necessary and sufficient conditions for spherical convexity.
The paper is structured as follows:
    In Section 2, we establish the general terminology used throughout the paper.
    In Section 3, we provide a brief overview of the main concepts related to spherical convexity and revisit a crucial result concerning the spherical convexity of smooth functions, which plays a paramount role in our study.
    Section 4 presents necessary, sufficient, and equivalent conditions for the spherical convexity of non-homogeneous quadratic functions. We specifically emphasize functions defined on the spherically convex set of unit vectors with positive coordinates, as well as functions corresponding to special classes of matrices—namely, positive matrices, Z-matrices, and diagonal matrices.
    Finally, in Section 5, we offer concluding remarks and propose challenging questions for future exploration.
2 Basic terminology
Let $n\geq 3$ be an integer and $\mathbb{R}^{n}\equiv\mathbb{R}^{n\times 1}$ be the vector space whose elements are column vectors of length $n$ and real entries.
Denote by $I\in\mathbb{R}^{n\times n}$ the identity matrix, that is, the matrix with elements $I_{ij}=\delta_{ij}$, where $\delta$ is the Kronecker symbol.
For any positive integer $k$ denote $[k]=\{1,\dots,k\}$. For all $i\in[n]$ denote by $e^{i}$ the column vector with all entries 00 except the $i$-th entry which is $1$. The basis $\{e^{1},\dots,e^{n}\}$ of $\mathbb{R}^{n}$ is called the standard basis. Denote $1=e^{1}+\dots+e^{n}$. For any vector $z\in\mathbb{R}^{n}$ let $z_{i}=z^{\top}e^{i}$, for all $i\in[n]$.
For any vector $d\in\mathbb{R}^{n}$ and symmetric matrix $A\in\mathbb{R}^{n\times n}$ denote $\operatorname{diag}(d):=\sum_{i=1}^{n}d_{i}e^{i}(e^{i})^{\top}\in\mathbb{R}^{n\times n}$, $\operatorname{diag}(A)=\sum_{i=1}^{n}\left\langle Ae^{i},e^{i}\right\rangle e^{i}\in\mathbb{R}^{n}$, $\operatorname{diag}^{2}(A)=\operatorname{diag}(\operatorname{diag}(A))\in\mathbb{R}^{n\times n}$ and $a_{ij}=\left\langle Ae^{i},e^{j}\right\rangle$ for all $i,j\in[n]$. Let $I:=\sum_{i=1}^{n}(e^{i})(e^{i})^{\top}$.
For any $x=(x_{1},\dots,x_{n})^{\top}$ and $y=(y_{1},\dots,y_{n})^{\top}$ in $\mathbb{R}^{n}$ denote $\left\langle x,y\right\rangle:=x^{\top}y$ the canonical inner product.
For any number $\alpha\in\mathbb{R}$ denote $\alpha_{+}=\max(\alpha,0)$, $\alpha_{-}=\max(-\alpha,0)$; for any vector $v=v_{1}e^{i}+\dots+v_{n}e^{n}=(v_{1},\dots,v_{n})^{\top}\in\mathbb{R}^{n}$ denote
$v_{+}=\left(v_{1}^{+},\dots,v_{n}^{+}\right)^{\top}\in\mathbb{R}^{n}_{+},\quad v_{-}=\left(v_{1}^{-},\dots,v_{n}^{-}\right)^{\top}\in\mathbb{R}^{n}_{+};$
and for any matrix $B\in\mathbb{R}^{k\times l}$ denote $B=(b_{ij})_{i\in[k],j\in[l]}$.
Let
$\mathbb{S}^{n}=\left\{x\in\mathbb{R}^{n}:\|x\|=1\right\}$
be the unit sphere, which we will shortly call just “sphere”.
With a slight abuse of notation we will use the same notation for the linear mappings of $\mathbb{R}^{n}$ and their matrices with respect to the standard basis.
A set $K\subseteq\mathbb{R}^{n}$ is called cone if $\lambda x\in K$, for all $x\in K$ and all $\lambda>0$. A cone $K\subseteq\mathbb{R}^{n}$ is called convex cone if is a if $x+y\in K$ for any $x,y\in K$. It follows that a cone $K\subseteq\mathbb{R}^{n}$ is a convex cone if and only if it is a convex set. A cone $K\subseteq\mathbb{R}^{n}$ is called pointed cone if there is no $x\in K\setminus\{0\}$ such that $-x\in K$. A cone $K\subseteq\mathbb{R}^{n}$ is called closed cone if it is a closed set. A cone $K\subseteq\mathbb{R}^{n}$ is called proper if it is a pointed closed convex cone with nonempty interior. The dual $K^{\perp}$ of a cone $K\subseteq\mathbb{R}^{n}$ is defined by
$K^{*}=\{x\in\mathbb{R}^{n}:\left\langle x,y\right\rangle\geq 0,\,\forall y\in K\}.$
The polar $K^{\perp}$ of $K$ is defined by
$K^{\perp}=\{x\in\mathbb{R}^{n}:\left\langle x,y\right\rangle\leq 0,\,\forall y\in K\}.$
3 Preliminaries on the sphere
The terminology and results about the sphere follow the ones in [9, 10]. Therefore, we will broadly describe the main concepts of spherical convexity and only repeat the formulas explicitly used later.
The geodesics of the sphere $\mathbb{S}^{n}$ are circles which are the intersection of two dimensional subspaces of $\mathbb{R}^{n}$ with the sphere.
The distance between any two points on the sphere is the length of the smallest geodesic arc joining the two points. The diameter of a subset of the sphere is the supremum of distances between all pair of points belonging to the sphere.
Any two points on the sphere of distance less than $\pi$ can be joined by two geodesic arcs and two diametrically opposed points by infinitely many. The smallest length geodesic arc joining two points is called minimal.
A set $\mathcal{C}\subseteq\mathbb{S}^{n}$ is called spherically convex if for any two points of $\mathcal{C}$ all minimal geodesic arcs joining the two points are contained in $\mathcal{C}$. A spherically convex set is called proper, if it is nonempty and it is not the whole sphere. This definition implicitly yields that the diameter of any proper spherically convex set is less than $\pi$ and thus the minimal geodesic arc joining any two points of the set is unique. The proper spherically convex sets are intersections of the sphere with pointed convex cones of $\mathbb{R}^{n}$.
A real valued function defined on a spherically convex set is called spherically convex if its composition with any minimal geodesic segment belonging to the set is a convex function defined on an interval. This definition implies that the only spherically convex functions on the whole sphere are the constant ones.
Denote by $D$ and $D^{2}$, the Euclidean gradient and Euclidean Hessian, respectively. The following proposition, proved in [13], will be paramount for our investigations.
Proposition 1.
Let $K\subseteq R^{n}$ be a proper cone, $\mathcal{C}=\mathbb{S}^{n}\cap\operatorname{int}(K)$ and $f:C\to\mathbb{R}$ a smooth function. Then, the following statements are equivalent:
  1. (i)
    $f$ is spherically convex;
  2. (ii)
    $\left\langle Df(x)-Df(y),x-y\right\rangle+(\left\langle x,y\right\rangle-1)[\left\langle Df(x),x\right\rangle+\left\langle Df(y),y\right\rangle]\geq 0$, for all $x,y\in\mathcal{C}$;
  3. (iii)
    $\left\langle D^{2}f(y)x,x\right\rangle-\left\langle Df(y),y\right\rangle\geq 0$, for all $y\in\mathcal{C}$, $x\in\mathbb{S}^{n}$ with $\left\langle x,y\right\rangle=0$.
4 Main results
Throughout this paper we restrict ourselves to proper cones $K\subseteq\mathbb{R}^{n}$, that is to pointed closed convex cones with nonempty interior. Let $c\in\mathbb{R}$, $A\in\mathbb{R}^{n\times n}$ be a symmetric matrix and $b\in\mathbb{R}^{n}$. From now on $f$ will always denote a function $f:\mathbb{S}^{n}\cap\operatorname{int}(K)\to\mathbb{R}$ defined by
$f(x)=\left\langle Ax,x\right\rangle+\left\langle b,x\right\rangle+c.$
Sometimes, to emphasize the dependence of $f$ on $A$, $b$, $c$, we will use the more specific notation $f=f_{A,b,c}$. Denote $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ the smallest and largest eigenvalues of $A$, respectively.
The following Proposition is a direct consequence of Proposition 1.
Proposition 2.
The following statements are equivalent:
  1. (i)
    $f$ is spherically convex;
  2. (ii)
    $\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle\geq\frac{1}{2}\left\langle b,v\right\rangle,$ (1)
    for all $u\in\mathbb{S}^{n}$ and all $v\in\mathbb{S}^{n}\cap K$ with $\left\langle u,v\right\rangle=0$.
  3. (iii)
    $4\left\langle Ax,y\right\rangle\leq 2\left\langle x,y\right\rangle\left(\left\langle Ax,x\right\rangle+\left\langle Ay,y\right\rangle\right)+(\left\langle x,y\right\rangle-1)\left\langle b,x+y\right\rangle,$ (2)
    for all $x,y\in\mathbb{S}^{n}\cap K$.
The following lemma demonstrates that the spherical convexity of $f$ remains unchanged if any constant multiple of the identity matrix is added to $A$.
Lemma 1.
Let $K$ be a proper cone, $A=A^{\top}\in\mathbb{R}^{n\times n}$, $b\in\mathbb{R}^{n}$ and $c,\lambda\in\mathbb{R}$. Then, $f_{A,b,c}$ is spherically convex if and only if $f_{A-\lambda I,b,c}$ is spherically convex.
Proof.
For any $u,v\in\mathbb{S}^{n}$ with $v\in K$ and $\left\langle u,v\right\rangle=0$ denote
$E(A,b,c,u,v):=2\left\langle Au,u\right\rangle-2\left\langle Av,v\right\rangle-\left\langle b,v\right\rangle.$
Then, according to (1), $f_{A,b,c}$ is spherically convex if and only if $E(A,b,c,u,v)\geq 0$ for all $u,v\in\mathbb{S}^{n}$ with $v\in K$ and $\left\langle u,v\right\rangle=0$. It is easy to check that $E(A-\lambda I,b,c,u,v)=E(A,b,c,u,v)$. Hence, the result readily follows. ∎
The following theorem lists necessary conditions, sufficient conditions, and equivalent conditions for $f$ to be spherically convex. The necessary conditions (ii)-(v) can serve as negative certificates for the spherical convexity of $f$. If any of these conditions do not hold, then $f$ is not spherically convex. The sufficient condition (vi) provides a positive certificate for the spherical convexity of $f$. If this condition holds, then $f$ is spherically convex. Moreover, this condition demonstrates that for any symmetric matrix $A\in\mathbb{R}^{n\times n}$, there are infinitely many $b$ such that $f$ is spherically convex, making the class of spherically convex non-homogeneous quadratic functions much larger than the class of (Euclidean) convex non-homogeneous quadratic functions. Note that the vector $b$ has no influence on the (Euclidean) convexity of $f$, but it is crucial for its spherical convexity. Even more interestingly, if $K=\mathbb{R}^{n}_{+}$, then this sharply contrasts with the fact that the class of spherically convex homogeneous functions is trivial, that is, the function $f$ is spherically convex for $b=0$ if and only if it is constant [13].
Theorem 1.
The following statements hold.
  1. (i)
    Suppose that $A=0$. Then, $f$ is spherically convex if and only if $b\in K^{\perp}$.
  2. (ii)
    Suppose that $K^{*}\subseteq K$ and $f$ is sperically convex. Then,
    $\left\langle Ax,y\right\rangle\leq-\frac{\sqrt{2}}{8}\left\langle b,x+y\right\rangle,$ (3)
    for all $x,y\in\mathbb{S}^{n}\cap K$ with $\left\langle x,y\right\rangle=0$. In particular, if $K=\mathbb{R}^{n}_{+}$, then
    $a_{ij}\leq-\frac{\sqrt{2}}{8}(b_{i}+b_{j}),$ (4)
    for all $i,j\in[n]$ with $i\neq j$.
  3. (iii)
    Suppose that $f$ is spherically convex. Then, $\left\langle b,x+y\right\rangle\leq 0$, for all $x,y\in\mathbb{S}^{n}\cap K$ with $\left\langle x,y\right\rangle=0$. In particular, if +$\mathbb{R}^{n}_{+}\subseteq K$ and $f$ is spherically convex, then
    $b_{i}+b_{j}\leq 0,$ (5)
    for all $i,j\in[n]$ with $i\neq j$, and
    $\min\left\{a_{ii}:i\in[n]\right\}\geq\max\left\{b_{j}+a_{jj}:j\in[n]\setminus\operatorname{argmin}\left\{a_{ii}:i\in[n]\right\}\right\}.$ (6)
  4. (iv)
    Suppose that $f$ is spherically convex. Then,
    $\left\langle b,x+y\right\rangle\leq-4\sqrt{2}\left\langle Ax,y\right\rangle^{+},$ (7)
    for all $x,y\in\mathbb{S}^{n}\cap K$ with $\left\langle x,y\right\rangle=0$. In particular, if $K=\mathbb{R}^{n}_{+}$, then
    $b_{i}+b_{j}\leq-4\sqrt{2}a_{ij}^{+},$ (8)
    for all $i,j\in[n]$ with $i\neq j$.
  5. (v)
    If $K=\mathbb{R}^{n}_{+}$, $A$ has positive entries and $f$ is spherically convex, then
    $\|b_{-}\|\geq 2\left[\lambda_{\max}(A)-\lambda_{\min}(A)\right].$ (9)
  6. (vi)
    If $K\subseteq\mathbb{R}^{n}_{+}$ and
    $b_{i}\leq 2\sqrt{n}\left[\lambda_{\min}(A)-\lambda_{\max}(A)\right],$
    for all $i\in[n]$, then $f$ is spherically convex.
Proof.
  1. (i)
    Suppose that $A=0$. Then, by using inequality (1), it follows that $f$ is spherically convex if and only if $\left\langle b,v\right\rangle\leq 0$ for all $v\in\mathbb{S}^{n}\cap K$. Therefore, $f$ is spherically convex if and only if $b\in K^{\perp}$.
  2. (ii)
    Inequality (3) is a consequence of inequality (1) with $u=\left(\sqrt{2}/2\right)(x-y)$ and $v=\left(\sqrt{2}/2\right)(x+y)$ because $v\in K$ and it can be easily verified that $\|u\|=\|v\|=1$, $\left\langle u,v\right\rangle=0$. In particular, if $K=\mathbb{R}^{n}_{+}$, then inequality (4) follows from (3) by taking $x=e^{i}$ and $y=e^{j}$.
  3. (iii)
    Suppose that $f$ is spherically convex and let any $x,y\in\mathbb{S}^{n}\cap K$ with $\left\langle x,y\right\rangle=0$. Then, the first inequality of the statement follows by adding the inequalities
    $\left\langle Ax,x\right\rangle-\left\langle Ay,y\right\rangle\geq\frac{1}{2}\left\langle b,y\right\rangle,\quad\left\langle Ay,y\right\rangle-\left\langle Ax,x\right\rangle\geq\frac{1}{2}\left\langle b,x\right\rangle,$
    which are consequences of inequality (1).
    The next inequality of the statement follows by taking any $i,j\in[n]$ with $i\neq j$ and $x=e^{i}$, $y=e^{j}$. On the other hand, inequality (1) implies
    $2(a_{ii}-a_{jj})\geq b_{j}.$ (10)
    Inequality (6) is a straightforward consequence of inequality (10).
  4. (iv)
    It straightforwardly follows from inequalities (3), (4), (7) and (5).
  5. (v)
    By using the Cauchy inequality we have
    $2\left[\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle\right]\geq\left\langle b,v\right\rangle=\left\langle b_{+}-b_{-},v\right\rangle\geq-\left\langle b_{-},v\right\rangle\geq-\|b_{-}\|.$ (11)
    Let $u,v$ be unit eigenvectors of $A$ corresponding to $\lambda_{\min}(A),\lambda_{\max}(A)$, respectively. From the Perron-Frobenius theorem [30, 15] it follows that we can choose $v\in\mathbb{\mathbb{R}}^{n}_{+}=K$, hence our choice of $u,v$ is feasible. Then, (11) implies (9).
  6. (vi)
    Since $v\in S\cap K$, it follows that there exists an $i_{0}\in[n]$ with $v_{i_{0}}\geq 1/\sqrt{n}$. Indeed, otherwise $0\leq v_{i}<1/\sqrt{n}$, for all $i\in[n]$ implies $\|v\|<1$, which contradicts $v\in\mathbb{S}^{n}$. We have
    $-b_{i}\geq 2\sqrt{n}\left[\lambda_{\max}(A)-\lambda_{\min}(A)\right]\geq 0,$
    for all $i\in[n]$, which implies
    $\frac{1}{2}\left\langle-b,v\right\rangle\geq\frac{1}{2}(-b_{i_{0}})v_{i_{0}}\geq\lambda_{\max}(A)-\lambda_{\min}(A)\geq\left\langle Av,v\right\rangle-\left\langle Au,u\right\rangle,$
    or equivalently
    $\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle\geq\frac{1}{2}\left\langle b,v\right\rangle.$
    Hence, inequality (1) implies that $f$ is spherically convex.
The next theorem presents a characterization of spherically convex functions. Although this characterization depends on a single vector $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$, it also involves an implicit term that resembles a scalar derivative, similar to those defined in [21] (see [28] for the earliest definition of the scalar derivative).
Theorem 2.
The function $f$ is spherically convex, if and only if
$\frac{1}{2}\left\langle b,x\right\rangle\leq\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}-\left\langle Ax,x\right\rangle,$
for all $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$.
Proof.
Suppose that $f$ is spherically convex. Let $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$ and $y\in\mathbb{S}^{n}\cap K$, such that $y\neq x$. Then, $01-\left\langle x,y\right\rangle>$ 0. Hence, according to (2) we have
$\left\langle b,x+y\right\rangle\leq\frac{4\left\langle x,y\right\rangle\left(\left\langle Ax,x\right\rangle+\left\langle Ay,y\right\rangle\right)-8\left\langle Ax,y\right\rangle}{2-2\left\langle x,y\right\rangle}.$ (12)
Let $t=\|y-x\|$ and $u=(1/t)(y-x)$. Then, $y=x+tu$ and $\|y\|=\|x\|=\|u\|=1$ imply
$2-2\left\langle x,y\right\rangle=2-2\left\langle x,x+tu\right\rangle=-2t\left\langle x,u\right\rangle=-\|x+tu\|^{2}+\|x\|^{2}+t^{2}\|u\|^{2}=t^{2}.$
Hence, (12) becomes
$2\left\langle b,x\right\rangle\leq\frac{\left(4-2t^{2}\|u\|^{2}\right)\left(2\left\langle Ax,x\right\rangle+2t\left\langle Ax,u\right\rangle+t^{2}\left\langle Au,u\right\rangle\right)-8\left\langle Ax,x\right\rangle-8t\left\langle Ax,u\right\rangle}{t^{2}}$
$=4\left\langle Au,u\right\rangle-4\left\langle Ax,x\right\rangle-4t\left\langle Ax,u\right\rangle-2t^{2}\left\langle Au,u\right\rangle-t\left\langle b,u\right\rangle,$
which implies
$2\left\langle b,x\right\rangle\leq\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\left(4\left\langle Au,u\right\rangle-4\left\langle Ax,x\right\rangle-4t\left\langle Ax,u\right\rangle-2t^{2}\left\langle Au,u\right\rangle-t\left\langle b,u\right\rangle\right).$
Hence,
$\frac{1}{2}\left\langle b,x\right\rangle\leq\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}-\left\langle Ax,x\right\rangle.$
Conversely, suppose that
$\frac{1}{2}\left\langle b,x\right\rangle\leq\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}-\left\langle Ax,x\right\rangle,$ (13)
for all $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$. Let $y=(\cos t)x+(\sin t)v$, where $t\in(0,2\pi)$ and $v\in\mathbb{S}^{n}\cap K$, $\left\langle v,x\right\rangle=0$. Then, $y\in\mathbb{S}^{n}\cap K$ and $y\neq x$. We have
$\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}=\frac{(\cos t-1)^{2}\left\langle Ax,x\right\rangle+(\sin(2t)-2\sin t)\left\langle Av,x\right\rangle+\sin^{2}t\left\langle Av,v\right\rangle}{2-2\cos t}$
Tending with $t$ to zero and using the L’Hopital rule twice we obtain
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}\leq\left\langle Av,v\right\rangle.$
Hence, (13) implies
$\frac{1}{2}\left\langle b,x\right\rangle\leq\left\langle Av,v\right\rangle-\left\langle Ax,x\right\rangle$
for all $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$ and any $v\in\mathbb{S}^{n}$ with $v\perp x$. By taking a limit and using continuity, we can assume that the same inequality holds for all $x\in\mathbb{S}^{n}\cap K$. Hence, (1) holds and therefore $f$ is spherically convex. ∎
Although the proofs of Theorem 2 and Proposition 3 suggest that Theorem 2 is merely a reformulation of Proposition 1 (at least when the matrix $A$ is positive definite), it leads to the following Corollary, which offers a new type of characterization of spherically convex non-homogeneous quadratic functions. This characterization depends only on a unit vector in the cone and an arbitrary unit vector that is not related to it. Previously, our characterizations involving two unit vectors either required that both vectors be in $K$ or that one unit vector be in $K$ and the other be perpendicular to it.
Corollary 1.
The function $f$ is spherically convex if and only if
$\frac{1}{2}\left\langle b,x\right\rangle\leq\frac{\left\langle Av,v\right\rangle-2\left\langle v,x\right\rangle\left\langle Av,x\right\rangle+(2\left\langle v,x\right\rangle^{2}-1)\left\langle Ax,x\right\rangle}{1-\left\langle v,x\right\rangle^{2}},$ (14)
for all $x\in\mathbb{S}^{n}\cap K$ and all unit vectors $v$ with $v\notin\{x,-x\}$.
Proof.
Let $f$ be spherically converse. First, suppose that $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$. Let $v$ be a unit vector such that $v\notin\{x,-x\}$. and $y=(x+tv)/\|x+tv\|$, where $t>0$ is such that $\|x+tv\|\neq 0$. If $t>0$ is sufficiently small then $\|x+tv\|\neq 0$ and $y\in\mathbb{S}^{n}\cap K$. Moreover $y\neq x$. First, let us calculate $\psi^{\prime}(0)$, where $\psi(t)=\|x+tv\|$. We have
$\frac{\psi(t)-\psi(0)}{t}=\frac{\|x+tv\|-1}{t}=\frac{(\|x+tv\|-1)(\|x+tv\|+1)}{t(\|x+tv\|+1}=\frac{\left\langle x+tv,x+tv\right\rangle-1}{t(\|x+tv\|+1}$
$=\frac{2\left\langle x,v\right\rangle+t\|v\|^{2}}{\|x+tv\|+1}$
Letting $t$ tend to zero in the above formula, we obtain
$\frac{d}{dt}|_{t=0}\|x+tv\|=\left\langle v,x\right\rangle.$
Multiplying the numerator and denominator of $\left\langle Ay-Ax,y-x\right\rangle/\|y-x\|^{2}$ by $\|x+tv\|^{2}$, we obtain
$\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}=\frac{\left\langle Ax+tAv-\|x+tv\|Ax,x+tv-\|x+tv\|x\right\rangle}{\left\langle x+tv-\|x+tv\|x,x+tv-\|x+tv\|x\right\rangle}$
Tending with $t$ to zero and using the L’Hopital rule twice, we obtain
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}\leq\frac{\left\langle Av,v\right\rangle-2\left\langle v,x\right\rangle\left\langle Av,x\right\rangle+\left\langle v,x\right\rangle^{2}\left\langle Ax,x\right\rangle}{1-\left\langle v,x\right\rangle^{2}}.$
Hence, Theorem 2 implies (14). If $v\in\{x,-x\}$, then let $v^{n}$ be unit vectors such that $v^{n}\to v$ if $n\to+\infty$. Writing inequality (14) with $v^{n}$ replacing $v$ and tending with $n$ to infinity, we obtain inequality (14). If $x$ is on the the intersection of $\mathbb{S}^{n}$ with the boundary of $K$, then there exists $x^{k}\in\mathbb{S}^{n}\cap\operatorname{int}(K)$ such that $\lim_{k\to\infty}x^{k}=x$. Hence,
$\frac{1}{2}\left\langle b,x^{k}\right\rangle\leq\frac{\left\langle Av,v\right\rangle-2\left\langle v,x^{k}\right\rangle\left\langle Av,x^{k}\right\rangle+(2\left\langle v,x^{k}\right\rangle^{2}-1)\left\langle Ax^{k},x^{k}\right\rangle}{1-\left\langle v,x^{k}\right\rangle^{2}},$
Tending with $k$ to infinity in the last inequality we obtain (14). Conversely suppose that
$\frac{1}{2}\left\langle b,x\right\rangle\leq\frac{\left\langle Av,v\right\rangle-2\left\langle v,x\right\rangle\left\langle Av,x\right\rangle+(2\left\langle v,x\right\rangle^{2}-1)\left\langle Ax,x\right\rangle}{1-\left\langle v,x\right\rangle^{2}},$
for all $x\in\mathbb{S}^{n}\cap K$ and all unit vectors $v$ with $v\notin\{x,-x\}$. Let $v\in\mathbb{S}^{n}$ such that $\left\langle v,x\right\rangle=0$. Then, inequality (1) holds and therefore $f$ is spherically convex. ∎
Without entering in details, we remark that the inferior limit in Theorem 2 is a variant of the scalar derivative defined by S. Z. Németh in [28] (for other variants see [21]). In the next proposition we show how to determine this inferior limit in the case when $A$ is positive definite. Note that, by Lemma 1, the positive definiteness of $A$ is not a restriction on the spherical convexity of $f=f_{A,b,c}$ because the spherical convexity of $f$ is equivalent to the spherical convexity of $f_{A+\alpha I,b,c}$ and we can take $\alpha$ sufficiently large for $A+\alpha I$ to be spherically convex.
Proposition 3.
Suppose that $A$ is positive definite and let $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$. Then,
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}=\min_{\begin{subarray}{c}u\in\mathbb{S}^{n}\\ \left\langle u,x\right\rangle=0\end{subarray}}\left\langle Au,u\right\rangle=\lambda_{\min}(P_{x}AP_{x}+\lambda Q_{x}),$ (15)
where $P_{x}=I-xx^{\top}$, $Q_{x}=xx^{\top}$ and $\lambda=\left\langle Ar,r\right\rangle$, for any fixed $r\in\mathbb{S}^{n}$ with $\left\langle r,x\right\rangle=0$.
Proof.
First we justify the second inequality in (15). Let $x^{\perp}=\{u\in\mathbb{R}^{n}\,:\,\left\langle u,x\right\rangle=0\}$. Please note the following properties of $P_{x}$ and $Q_{x}$:
$\begin{array}[]{r}P_{x}+Q_{x}=I,\quad P_{x}^{2}=P_{x}=P_{x}^{\top},\quad Q_{x}^{2}=Q_{x}=Q_{x}^{\top},\quad P_{x}Q_{x}=0,\quad P_{x}x=0,\quad Q_{x}x=x,\\ w\in\mathbb{R}^{n}\implies P_{x}w=w-\left\langle w,x\right\rangle x\in x^{\perp},\quad z\in x^{\perp}\implies P_{x}z=z\end{array}$ (16)
Formula (16)7 implies that $P_{x}AP_{x}x^{\perp}\subseteq x^{\perp}$. Let any $z\in x^{\perp}$. Since $A$ is positive definite, $A^{-1}$ is also positive definite. Therefore, $\left\langle A^{-1}x,x\right\rangle>0$ and the vector
$v:=A^{-1}z-\frac{\left\langle A^{-1}z,x\right\rangle}{\left\langle A^{-1}x,x\right\rangle}A^{-1}x\in x^{\perp}$
is well defined. By using (16)8,5, we obtain $P_{x}AP_{x}x^{\perp}\ni P_{x}AP_{x}v=P_{x}Av=z$. Hence, $x^{\perp}\subseteq P_{x}AP_{x}x^{\perp}$. In conclusion $P_{x}AP_{x}x^{\perp}=x^{\perp}$ and therefore $x^{\perp}$ is an invariant hyperplane of $P_{x}AP_{x}$. With a slight abuse of notation, we will identify the linear operators of $\mathbb{R}^{n}$ with their matrices with respect to the standard canonical basis. Then, we have
$\lambda_{\min}(P_{x}AP_{x}|_{x^{\perp}})=\min_{\begin{subarray}{c}u\in\mathbb{S}^{n}\\ \left\langle u,x\right\rangle=0\end{subarray}}\left\langle Au,u\right\rangle.$
To justify the second inequality in (15), it remains to show that
$\lambda_{\min}(P_{x}AP_{x}+\lambda Q_{x})=\lambda_{\min}(P_{x}AP_{x}|_{x^{\perp}}).$ (17)
Let $w\in\mathbb{R}^{n}\setminus\{0\}$ be an eigenvector of $P_{x}AP_{x}+\lambda Q_{x}$ with $P_{x}w\neq 0$ and corresponding eigenvalue $\mu$. Multiplying the left and right hand sides of
$(P_{x}AP_{x}+\lambda Q_{x})w=\mu w$ (18)
by $P_{x}$ and using (16)2,4,7, we obtain $P_{x}AP_{x}P_{x}w=\mu P_{x}w$, which implies that $\mu$ is an eigenvalue of $P_{x}AP_{x}|_{x^{\perp}}$ with corresponding eigenvector $P_{x}w\in x^{\perp}$. Hence,
$\mu\geq\lambda_{\min}(P_{x}AP_{x}|_{x^{\perp}}).$ (19)
Next, let $w\in\mathbb{R}^{n}\setminus\{0\}$ be an eigenvector of $P_{x}AP_{x}+\lambda Q_{x}$ with $P_{x}w=0$ and the corresponding eigenvalue $\mu$. Since $P_{x}w=0$, it follows form (16)7 that $w=\left\langle w,x\right\rangle x$. Hence, equations (16)5,6 and (18) imply
$\mu\left\langle w,x\right\rangle=(P_{x}AP_{x}+\lambda Q_{x})\left\langle w,x\right\rangle=\lambda\left\langle w,x\right\rangle.$ (20)
Note that $\left\langle w,x\right\rangle\neq 0$ because otherwise $P_{x}w=w-\left\langle w,x\right\rangle x=w\neq 0$ contradicts our assumption $P_{x}w=0$. Thus, (20), $v\in S\cap x^{\perp}$ and (16)2,8 implies
$\mu=\lambda=\left\langle Ar,r\right\rangle=\left\langle AP_{x}r,P_{x}r\right\rangle=\left\langle P_{x}AP_{x}r,r\right\rangle\geq\lambda_{\min}(P_{x}AP_{x}|_{x^{\perp}}).$
Hence,
$\mu\geq\lambda_{\min}(P_{x}AP_{x}|_{x^{\perp}}).$ (21)
Inequalities (19) and (21) imply that
$\lambda_{\min}(P_{x}AP_{x}+\lambda Q_{x})\geq\lambda_{\min}(P_{x}AP_{x}|_{x^{\perp}}).$ (22)
Conversely, let $a\in x^{{}^{\perp}}$ be an eigenvector of $P_{x}AP_{x}|_{x^{\perp}}$ with corresponding eigenvalue $\mu$. Then $P_{x}AP_{x}a=\mu a$ and (16)8,4 imply
$(P_{x}AP_{x}+\lambda Q_{x})a=(P_{x}AP_{x})a+\lambda Q_{x}a=\mu a+\lambda Q_{x}P_{x}a=\mu a,$
which implies that $\mu$ is an eigenvalue of $P_{x}AP_{x}+\lambda Q_{x}$ and thus $\mu\geq\lambda_{\min}(P_{x}AP_{x}+\lambda Q_{x})$. In particular,
$\lambda_{\min}(P_{x}AP_{x}|_{x^{\perp}})\geq\lambda_{\min}(P_{x}AP_{x}+\lambda Q_{x}).$ (23)
Inequalities (22) and (23) imply inequality (17) and in conclusion (17)2 holds.
To show the first equality in (15) please note that we have already shown in the proof of Theorem 2 that
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}\leq\left\langle Au,u\right\rangle,$
for all $u\in\mathbb{S}^{n}$ with $u\perp x$, which implies
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}\leq\min_{\begin{subarray}{c}u\in\mathbb{S}^{n}\\ \left\langle u,x\right\rangle=0\end{subarray}}\left\langle Au,u\right\rangle.$
It remains to show that
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}\geq\min_{\begin{subarray}{c}u\in\mathbb{S}^{n}\\ \left\langle u,x\right\rangle=0\end{subarray}}\left\langle Au,u\right\rangle.$ (24)
Let any $\varepsilon>0$ and $y^{k}\in\mathbb{S}^{n}\cap K$ be a sequence such that $y^{n}\neq x$ for all positive integer $k$, $\lim_{k\to\infty}y^{k}=x$ and the sequence
$\frac{\left\langle Ay^{k}-Ax,y^{k}-x\right\rangle}{\|y^{k}-x\|^{2}}=\left\langle A\left(\frac{y^{k}-x}{\|y^{k}-x\|}\right),\frac{y^{k}-x}{\|y^{k}-x\|}\right\rangle=\left\langle Az^{k},z^{k}\right\rangle$ (25)
is convergent to
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}-\varepsilon,$ (26)
where $z^{k}:=(y^{k}-x)/\|y^{k}-x\|$. Since $z^{k}\in\mathbb{S}^{n}$ and $\mathbb{S}^{n}$ is compact, it follows that there exists a convergent subsequence $z^{k_{\ell}}$ of $z^{k}$ such that
$v:=\lim_{\ell\to\infty}z^{k_{\ell}}\in\mathbb{S}^{n}.$ (27)
We have
$\lim_{\ell\to\infty}\left\langle Az^{k_{\ell}},z^{k_{\ell}}\right\rangle=\left\langle Av,v\right\rangle.$ (28)
On the other hand we have
$\left\langle y^{k_{\ell}}+x,z^{k_{\ell}}\right\rangle=\left\langle y^{k_{\ell}}+x,\frac{y^{k_{\ell}}-x}{\|y^{k_{\ell}}-x\|}\right\rangle=\frac{\|y^{k_{\ell}}\|^{2}-\|x\|^{2}}{\|y^{k_{\ell}}-x\|}=0$
Hence,
$\left\langle v,x\right\rangle=\frac{1}{2}\lim_{\ell^{\top}\infty}\left\langle y^{k_{\ell}}+x,z^{k_{\ell}}\right\rangle=0$ (29)
Formula (25) with $k_{\ell}$ replacing $k$, (26),(27), (28) and (29) imply
$\liminf_{\mathclap{\begin{subarray}{c}y\to x\\ \|y\|=1\\ y\in K\end{subarray}}}\frac{\left\langle Ay-Ax,y-x\right\rangle}{\|y-x\|^{2}}-\varepsilon=\left\langle Av,v\right\rangle\geq\min_{\begin{subarray}{c}u\in\mathbb{S}^{n}\\ \left\langle u,x\right\rangle=0\end{subarray}}\left\langle Au,u\right\rangle,$
for all $\varepsilon>0$, which implies (24) and in conclusion (24) holds. ∎
Proposition 3 and Theorem 2 (or equation (1)) imply the following theorem for $x\in\mathbb{S}^{n}\cap\operatorname{int}(K)$. To extend the theorem to $x\in\mathbb{S}^{n}\cap K$, one needs to approach $x$ with a sequence of points in $\mathbb{S}^{n}\cap\operatorname{int}(K)$ and use continuity.
Theorem 3.
Suppose that $A$ is positive definite. The function $f$ is spherically convex if and only if
$\frac{1}{2}\left\langle b,x\right\rangle\leq\lambda_{\min}(P_{x}AP_{x}+\lambda Q_{x})-\left\langle Ax,x\right\rangle,$
for all $x\in\mathbb{S}^{n}\cap K$, where $P_{x}=I-xx^{\top}$, $Q_{x}=xx^{\top}$ and $\lambda=\left\langle Ar,r\right\rangle$, for some $r\in\mathbb{S}^{n}$ with $\left\langle r,x\right\rangle=0$.
The following proposition is related to item (iii) of Theorem 1. It shows that the condition $\langle b,u+v\rangle\leq 0$ for arbitrary perpendicular unit vectors with nonnegative components is no more general than the same condition with two different vectors from the standard basis.
Proposition 4.
Let $K=\mathbb{R}^{n}_{+}$, $b\in\mathbb{R}^{n}$ and $u,v\in\mathbb{S}^{n}\cap\mathbb{R}^{n}_{+}$ such that $\left\langle u,v\right\rangle=0$. If $b_{i}+b_{j}\leq 0$, for all $i,j\in[n]$ with $i\neq j$, then $\left\langle b,u+v\right\rangle\leq 0$.
Proof.
If $b_{i}\leq 0$, for all $i\in[n]$, then the statement is trivial. Suppose that $b$ has a positive coordinate. By permutating the coordinates, we can suppose without loss of generality that $b_{1}>0$, $b_{i}\leq 0$, for all $i\in[n]\setminus\{1\}$, $u=\sum_{i=1}^{k}u_{i}e^{i}$ and $v=\sum_{j=k+1}^{n}v_{j}e^{j}$. Then,
$\langle b,u+v\rangle=b_{1}u_{1}+\sum_{i=2}^{k}b_{i}u_{i}+\sum_{j=k+1}^{n}b_{j}v_{j}\leq b_{1}u_{1}+\sum_{j=k+1}^{n}b_{j}v_{j}\leq b_{1}+\sum_{j=k+1}^{n}b_{j}v_{j}$
$\leq b_{1}+\max(b_{k+1},\dots,b_{n})\sum_{j=k+1}^{n}v_{j}\leq b_{1}+\max(b_{k+1},\dots,b_{n})\sum_{j=k+1}^{n}v_{j}^{2}=b_{1}+\max(b_{k+1},\dots,b_{n})$
$\leq 0.$
Note that, according to item (iii) of Theorem 1, it is easy to prove that if $f$ is spherically convex, then at most one component of $b$ can be positive. The following proposition demonstrates that if $f$ is spherically convex, then the components of $b$ corresponding to the non-minimal elements of the diagonal of $A$ are nonpositive. Additionally, it provides an upper bound for each component of $b$ corresponding to a minimal element of the diagonal of $A$. Some specific cases are also presented.
Proposition 5.
Let $K=\mathbb{R}^{n}_{+}$ and any $m\in\operatorname{argmin}\{a_{\ell\ell}:\ell\in[n]\}$. If $f$ is spherically convex, then $b_{i}\leq 0$, for all $i\in[n]\setminus\{m\}$ and
$b_{m}\leq-\max\left\{b_{i}+4a_{mi}^{+}:i\in[n]\setminus\{m\}\right\}.$ (30)
In particular, if $a_{11}=\dots=a_{nn}$ or $4a_{mi}^{+}\geq-b_{i}$, for some $m\in\operatorname{argmin}\{a_{\ell\ell}:\ell\in[n]\}$ and some $i\in[n]\setminus\{m\}$, then $b_{i}\leq 0$ for all $i\in[n]$.
Proof.
The inequality $b_{i}\leq 0$ for any $i\in[n]\setminus\{m\}$ follows from inequality (1) with $u=e^{j}$ and $v=e^{i}$, for any $i,j\in[n]$ with $i\neq j$. Inequality (30) follows from inequality (8). ∎
The following proposition shows that for a spherically convex $f$, if we delete the row and column of $A$ corresponding to any of its minimal diagonal elements and impose that the eigenvalue of the resulting matrix is less than the corresponding diagonal element, then all components of $b$ are nonpositive.
Proposition 6.
Let $K=\mathbb{R}^{n}_{+}$, and for any $k\in[n]$ denote $A_{-k}=(a_{ij})_{i,j\in[n]\setminus\{k\}}$. If $\lambda_{\min}(A_{-m})\leq a_{mm}$, for all $m\in\operatorname{argmin}\{a_{\ell\ell}:\ell\in[n]\}$ and and $f$ is spherically convex, then $b_{i}\leq 0$, for all $i\in[n]$.
Proof.
Let any $m\in\operatorname{argmin}\{a_{\ell\ell}:\ell\in[n]\}$. Take $u=\operatorname{argmin}\{\left\langle Ax,x\right\rangle:x\in\mathbb{S}^{n},x^{\top}e^{m}=0\}$ and $v=e^{m}$. Then, $\left\langle Au,u\right\rangle=\lambda_{\min}(A_{-m})$. Hence, inequality (1) implies that
$0\geq\lambda_{\min}(A_{-m})-a_{mm}=\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle\geq\frac{1}{2}\left\langle b,v\right\rangle=b^{m}.$
The inequalities $b_{i}\leq 0$, for any $i\in[n]\setminus\{m\}$ follows from proposition 5. ∎
The following theorem provides sufficient conditions for $f$ to be spherically convex when $K$ is the nonnegative orthant, based on the copositivity of certain matrices dependent on $A$. Additionally, item (iii) demonstrates that the inequality in item (i) is implied by item (ii).
Theorem 4.
Let $K=\mathbb{R}^{n}_{+}$. Consider the following statements
  1. (i)
    The matrix $\operatorname{diag}^{2}(A)-A$ is copositive and $b_{i}\leq 2[\lambda_{\min}(A)-a_{ii}],$ for all $i\in[n]$.
  2. (ii)
    The matrix $2[\lambda_{\min}(A)I-A]-\operatorname{diag}(b)$ is copositive and $b_{i}\leq 0$, for all $i\in[n]$.
  3. (iii)
    The inequality $b_{i}\leq 2\left[\lambda_{\min}(A)-a_{ii}\right]$ holds for all $i\in[n]$.
  4. (iv)
    The function $f$ is sperically convex.
Then, we have the following implications: (ii)$\implies$(iii) and (i)$\implies$(ii)$\implies$(iv).
Proof.
Let $B:=\operatorname{diag}^{2}(A)-A$ and $C:=2[\lambda_{\min}(A)I-A]-\operatorname{diag}(b)$.
(ii)$\implies$(iii): Let any $i\in[n]$. Since $C$ is copositive and $e^{i}\in K$, we have $\left\langle Ce^{i},e^{i}\right\rangle=2[\lambda_{\min}(A)-a_{ii}]-b_{i}\geq 0$, which implies $b_{i}\leq 2[\lambda_{\min}(A)-a_{ii}].$
(i)$\implies$(ii): Since $b_{i}\leq 2[\lambda_{\min}(A)-a_{ii}],$ we have that $2\lambda_{\min}(A)I-2\operatorname{diag}^{2}(A)-\operatorname{diag}(b)$ is a diagonal matrix with nonnegative elements, and hence copositive. Thus, we have that $C=2B+2\lambda_{\min}(A)I-2\operatorname{diag}^{2}(A)-\operatorname{diag}(b)$ is copositive, because it is the sum of two copositive matrices. Since $e^{i}\in\mathbb{S}^{n}$, we also have $b_{i}\leq 2[\lambda_{\min}(A)-a_{ii}]=\min_{z\in\mathbb{S}^{n}}\left\langle Az,z\right\rangle-\left\langle Ae^{i},e^{i}\right\rangle\leq 0$.
(ii)$\implies$(iv): Let any $u,v\in\mathbb{S}^{n}$ such that $v\in K$. Denote by $v^{2}\in K^{\*}$ the vector of coordinates $(v^{i})^{2}$ and by $u^{2}\in K^{\*}$ the vector of coordinates $(u_{i})^{2}$. Since $v\in\mathbb{S}^{n}\cap K$, we have $v-v^{2}\in K=K^{\*}$. Thus, the copositivity of $C$, $-b\in K=K^{\*}$ and $v-v^{2}\in K^{\*}$ imply
$2[\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle]=2[\left\langle Au,u\right\rangle-\lambda_{\min}(A)]-2[\left\langle Av,v\right\rangle-\lambda_{\min}(A)]$
$=2[\left\langle Au,u\right\rangle-\min_{z\in\mathbb{S}^{n}}\left\langle Az,z\right\rangle]-2[\left\langle Av,v\right\rangle-\lambda_{\min}(A)]\geq\left\langle 2[\lambda_{\min}(A)I-A]v,v\right\rangle$
$=\left\langle Cv,v\right\rangle+\left\langle\operatorname{diag}(b)v,v\right\rangle\geq\left\langle\operatorname{diag}(b)v,v\right\rangle=\left\langle b,v^{2}\right\rangle=\left\langle-b,v-v^{2}\right\rangle+\left\langle b,v\right\rangle\geq\left\langle b,v\right\rangle.$
Hence, the inequality (1) holds and therefore $f$ is spherically convex
The following corollary provides a more specific certificate for the spherical convexity of $f$ when $A$ is a $Z$-matrix.
Corollary 2.
Let $K=\mathbb{R}^{n}_{+}$. Suppose that $A$ is a $Z$-matrix and
$b_{i}\leq 2\left[\lambda_{\min}(A)-a_{ii}\right],$
for all $i\in[n]$. Then, $f$ is spherically convex.
Proof.
It follows from Theorem 4 (i)$\implies$(iv) because $\operatorname{diag}^{2}(A)-A$ is a matrix with nonnegative elements and hence it is copositive. ∎
The next corollary is a specialization of Corollary 2 for diagonal matrices.
Corollary 3.
Let $K=\mathbb{R}^{n}_{+}$, $d\in\mathbb{R}^{n}$ and $A=\operatorname{diag}(d)$ be a diagonal matrix. If
$b_{i}\leq 2\left[\min(d_{1},\dots,d_{n})-d_{i}\right]$
for all $i\in[n]$, then $f$ is spherically convex.
Proof.
It follows from Corollary 2 because $A$ is a $Z$-matrix and $\lambda_{min}(A)=\min(d_{1},\dots,d_{n})$. ∎
The next theorem provides a necessary and sufficient condition for the spherical convexity of $f$ when $A$ contains at least two minimal diagonal elements.
Theorem 5.
Let $K=\mathbb{R}^{n}_{+}$, $d\in\mathbb{R}^{n}$ and $A=\operatorname{diag}(d)$ be a diagonal matrix such that $\exists i_{0},j_{0}\in\operatorname{argmin}\{d_{i}:i\in[n]\}$ with $i_{0}\neq j_{0}$. Then, $f$ is spherically convex if and only if
$b_{i}\leq 2\left[\min(d_{1},\dots,d_{n})-d_{i}\right]$
for all $i\in[n]$.
Proof.
Suppose that $f$ is spherically convex and let any $i\in[n]$. Choose any
$i_{0}\in\operatorname{argmin}\left\{d_{i}:i\in[n]\right\}$
with $i_{0}\neq i$. Let $u=e^{i_{0}}$ and $v=e^{i}$. Then inequality (1) implies $2\left(d_{i_{0}}-d_{i}\right)\geq b_{i}$. Hence,
$b_{i}\leq 2\left[\min(d_{1},\dots,d_{n})-d_{i}\right].$
The converse follows from Corollary 3. ∎
The following lemma has a crucial role in the proof of Lemma 3.
Lemma 2.
Let $a,u,v\in\mathbb{R}^{n}$ such that $\|a\|=\|u\|=\|v\|=1$ and $\left\langle u,v\right\rangle=0$. Then,
$\left\langle a,u\right\rangle^{2}+\left\langle a,v\right\rangle^{2}\leq 1,$
Proof.
Let $\theta$ be the angle between $u$ and $a$, and $\varphi$ be the angle between $v$ and $a$. If $a=\pm u$ or $a=\pm v$, then $\left\langle a,u\right\rangle^{2}+\left\langle a,v\right\rangle^{2}=1$. Suppose that $a\neq\pm u$ and $a\neq\pm v$. We have $u=(\cos\theta)a+(\sin\theta)b$ and $u=(\cos\varphi)a+(\sin\varphi)c$, where $\|b\|=\|c\|=1$ and $\left\langle a,b\right\rangle=\left\langle a,c\right\rangle=0$. The formula
$0=\left\langle u,v\right\rangle=\cos\theta\cos\varphi+\sin\theta\sin\varphi\left\langle b,c\right\rangle$
implies $\cot\theta\cot\varphi=-\left\langle b,c\right\rangle\in[-1,1].$ Hence $\cot^{2}\theta\cot^{2}\varphi\leq 1$, which yields
$\left\langle a,u\right\rangle^{2}+\left\langle a,v\right\rangle^{2}=\cos^{2}\theta+\cos^{2}\varphi=2-\sin^{2}\theta-\sin^{2}\varphi=2-\frac{1}{1+\cot^{2}\theta}-\frac{1}{1+\cot^{2}\varphi}\leq 1$
Due to Lemma 1, the following lemma is essentially not less general than Theorem 6. We present it here primarily because proving it is technically more convenient than directly proving Theorem 6.
Lemma 3.
Let $K=\mathbb{R}^{n}_{+}$, $d\in\mathbb{R}^{n}_{+}$, $\tau$ be a permuation of $[n]$ such that
$0=d_{\tau(1)}<d_{\tau(2)}=\dots=d_{\tau(n)}$ (31)
and $A=\operatorname{diag}(d)$. If $b_{\tau(1)}\leq 2d_{\tau(2)}$ and
$b_{i}\leq-2d_{\tau(2)}\sqrt{6\sqrt{3}-9},$
for all $i\in[n]\setminus\{\tau(1)\}$, then $f$ is spherically convex.
Proof.
Suppose that $b_{\tau(1)}\leq 2d_{\tau(2)}$ and $b_{i}\leq-2d_{\tau(2)}\sqrt{6\sqrt{3}-9}$, which is equivalent to $b_{i}\leq-2d_{i}-\alpha$, where
$\alpha:=2\left(-1+\sqrt{6\sqrt{3}-9}\right)d_{\tau(2)}\approx(0.36)d_{\tau(2)}>0,$
for all $i\in[n]\setminus\{\tau(1)\}$. If $b_{\tau(1)}\leq 0$, then the spherical convexity of $f$ follows form Corollary 3. Suppose that $b_{\tau(1)}>0$. Then, $v\in K$, $b_{\tau(1)}\leq 2d_{\tau(2)}$ and $b_{i}\leq-2d_{i}-\alpha$, for all $i\in[n]\setminus\{\tau(1)\}$ implies $b_{\tau(1)}v_{\tau(1)}\leq 2d_{\tau(2)}v_{\tau(1)}$, $\left\langle b,v\right\rangle-b_{\tau(1)}v_{\tau(1)}\leq-2\left\langle d,v\right\rangle-\alpha\left\langle 1,v\right\rangle+\alpha v_{\tau(1)}$. By summing up the last two inequalities, we obtain
$\left\langle b,v\right\rangle=b_{\tau(1)}v_{\tau(1)}+\left\langle b,v\right\rangle-b_{\tau(1)}v_{\tau(1)}\leq 2d_{\tau(2)}v_{\tau(1)}-2\left\langle d,v\right\rangle-\alpha\left\langle 1,v\right\rangle+\alpha v_{\tau(1)}.$ (32)
We also have
$2\left\langle d,v\right\rangle+\alpha\left\langle 1,v\right\rangle-\alpha v_{\tau(1)}\geq\left(2d_{\tau(2)}+\alpha\right)\left(\left\langle 1,v\right\rangle-v_{\tau(1)}\right)$
$\geq\left(2d_{\tau(2)}+\alpha\right)\sqrt{\left\langle 1,v^{2}\right\rangle-v_{\tau(1)}^{2}}=\left(2d_{\tau(2)}+\alpha\right)\sqrt{\|v\|^{2}-v_{\tau(1)}^{2}}=\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$ (33)
because $v_{i}\geq 0$ for all $i\in[n]$. By combining (32) and (4) and using the equations (31), $d_{\tau(1)}=0$, $1=\|u\|^{2}=\left\langle 1,u^{2}\right\rangle$, $1=\|v\|^{2}=\left\langle 1,v^{2}\right\rangle$, we get
$2\left\langle d,u^{2}\right\rangle-2\left\langle d,v^{2}\right\rangle-\left\langle b,v\right\rangle\geq 2\left\langle d,u^{2}\right\rangle-2\left\langle d,v^{2}\right\rangle-2d_{\tau(2)}v_{\tau(1)}+2\left\langle d,v\right\rangle+\alpha\left\langle 1,v\right\rangle-\alpha v_{\tau(1)}$
$\geq 2\left\langle d,u^{2}\right\rangle-2\left\langle d,v^{2}\right\rangle-2d_{\tau(2)}v_{\tau(1)}+\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$
$=2d_{\tau(2)}\left\langle 1,u^{2}-u^{2}_{\tau(1)}e^{\tau(1)}\right\rangle-2d_{\tau(2)}\left\langle 1,v^{2}-v^{2}_{\tau(1)}e^{\tau(1)}\right\rangle-2d_{\tau(2)}v_{\tau(1)}+\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$
$=2d_{\tau(2)}\left\langle 1,u^{2}\right\rangle-2d_{\tau(2)}u^{2}_{\tau(1)}-\left(2d_{\tau(2)}\left\langle 1,v^{2}\right\rangle-2d_{\tau(2)}v^{2}_{\tau(1)}\right)-2d_{\tau(2)}v_{\tau(1)}$
$+\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$
$=2d_{\tau(2)}\|u\|^{2}-2d_{\tau(2)}u^{2}_{\tau(1)}-\left(2d_{\tau(2)}\|v\|^{2}-2d_{\tau(2)}v^{2}_{\tau(1)}\right)-2d_{\tau(2)}v_{\tau(1)}$
$+\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$
$=2d_{\tau(2)}-2d_{\tau(2)}u^{2}_{\tau(1)}-\left(2d_{\tau(2)}-2d_{\tau(2)}v^{2}_{\tau(1)}\right)-2d_{\tau(2)}v_{\tau(1)}+\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$
$=2d_{\tau(2)}\left(v^{2}_{\tau(1)}-u^{2}_{\tau(1)}-v_{\tau(1)}\right)+\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$ (34)
From Lemma 2 with $a=e^{\tau(1)}$ we have $-u_{\tau(1)}^{2}\geq v_{\tau(1)}^{2}-1$, which combined with (34) imply
$2\left\langle d,u^{2}\right\rangle-2\left\langle d,v^{2}\right\rangle-\left\langle b,v\right\rangle\geq 2d_{\tau(2)}\left(2v^{2}_{\tau(1)}-v_{\tau(1)}-1\right)+\left(2d_{\tau(2)}+\alpha\right)\sqrt{1-v_{\tau(1)}^{2}}$
$=2d_{\tau(2)}\left(2x^{2}-x-1+\beta\sqrt{1-x^{2}}\right)\geq 0,$ (35)
where $x:=v_{\tau(1)}$ and $\beta:=(2d_{\tau(2)}+\alpha)/2d_{\tau(2)}=\sqrt{6\sqrt{3}-9}$, and the last inequality in (35) above can be checked by showing that $\beta$ is the maximal value of the function $\psi:\mathbb{R}\setminus\{-1,1\}\to\mathbb{R}$ defined by
$\psi(x)=\frac{-2x^{2}+x+1}{\sqrt{1-x^{2}}}$
in the interval $[0,1)$. Indeed, it can be easily verified that $\psi$ is concave and its unique stationary point is $x^{*}=(\sqrt{3}-1)/2\in[0,1)$ (which is its unique global maximiser) and $\psi(x^{*})=\beta$. Hence $\beta=\max_{x\in[0,1)}\psi(x)$ and the last inequality in (35) readily follows. ∎
The next theorem demonstrates that there exist spherically convex functions $f$ such that one component of $b$ is positive.
Theorem 6.
Let $K=\mathbb{R}^{n}_{+}$, $\tau$ be a permuation of $[n]$ such that
$d_{\tau(1)}<d_{\tau(2)}=\dots=d_{\tau(n)}$
and $A=\operatorname{diag}(d)$. If $b_{\tau(1)}\leq 2\left(d_{\tau(2)}-d_{\tau(1)}\right)$ and
$b_{i}\leq-2\left(d_{\tau(2)}-d_{\tau(1)}\right)\sqrt{6\sqrt{3}-9},$
for all $i\in[n]\setminus\{\tau(1)\}$, then $f$ is spherically convex.
Proof.
From Lemma 1 $f$ is spherically convex if and only if $f_{A-d_{\tau(1)}I,b,c}$ is spherically convex. Hence, the result readily follows. ∎
The next proposition can be used as a further negative certificate for spherically convex functions. Indeed, if $f$ does not satisfy the condition of the Proposition for some indices $i\neq j$, then $f$ is not spherically convex. Solving the corresponding minimization problem is just a matter of simple numerical analysis.
Proposition 7.
Let $K=\mathbb{R}^{n}_{+}$. If $f$ is sperically convex, then
$\min\left\{\left[(a_{ii}-a_{jj})\cos(2\theta)-2a_{ij}\sin(2\theta)-\sin(\theta)b_{i}-\cos(\theta)b_{j}\right]\,:\,\theta\in\left[0,\frac{\pi}{2}\right]\right\}\geq 0,$
for all $i,j\in[n]$ with $i\neq j$.
Proof.
Let any $\theta\in[0,\pi/2]$ and any $i,j\in[n]$ with $i\neq j$. Then, the result follows from (1) with $u=\cos{\theta}e^{i}-\sin{\theta}e^{j}\in\mathbb{S}^{n}$ and $v=\sin(\theta)e^{i}+\cos(\theta)e^{j}\in\mathbb{S}^{n}\cap K$ after some trigonometric manipulations. ∎
The next two lemmas are building blocks of Theorem 7. Although in the case $n>2$ Lemma 4 follows from Theorem 5, we give an explicit proof below which works for all $n$.
Lemma 4.
Let $K=\mathbb{R}^{n}_{+}$, $i\in[n]$ and $A=\pm e^{i}(e^{i})^{\top}$. Then, $f$ is spherically convex if and only if $b_{j}\leq-2\delta_{ij}$, for all $j\in[n]$.
Proof.
Suppose that $f$ is spherically convex. Let any $j,k\in[n]$ such that $j\neq k$ and $k\neq i$. If $A=e^{i}(e^{i})^{\top}$, then the inequality follows by using inequality (1) with $u=e^{k}$, $v=e^{j}$. If, $A=-e^{i}(e^{i})^{\top}$, then the inequality follows by using inequality (1) with $u=e^{j}$, $v=e^{k}$.
Conversely, suppose that $b_{j}\leq-2\delta_{ij}$, for all $j\in[n]$. Let any $u\in\mathbb{S}^{n}$ and any $v\in\mathbb{S}^{n}\cap K$ with $\left\langle u,v\right\rangle=0$. If $A=e^{i}(e^{i})^{\top}$, then we have
$\frac{1}{2}\left\langle b,v\right\rangle\leq\frac{1}{2}b_{i}v_{i}\leq-v_{i}\leq-v_{i}^{2}\leq u_{i}^{2}-v_{i}^{2}=\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle,$
which implies inequality (1). Therefore, $f$ is spherically convex. If $A=-e^{i}(e^{i})^{\top}$, then inequality (1) trivially holds whenever $u_{i}^{2}-v_{i}^{2}\geq 0$. Therefore, to show that (1) holds, we can assume without loss of generality that $v_{i}^{2}-u_{i}^{2}>0$. Hence, we have
$\frac{1}{2}\left\langle b,v\right\rangle\leq\frac{1}{2}b_{i}v_{i}\leq-v_{i}\leq v_{i}^{2}-u_{i}^{2}=\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle,$
which implies inequality (1). Thus, $f$ is spherically convex. ∎
Although only item (ii) of the next lemma is used in the proof of Theorem 7, we have included item (i) as well to make the result more complete.
Lemma 5.
Let $K=\mathbb{R}^{n}_{+}$, $i,j\in[n]$, $i\neq j$ and $A=\pm e^{i}(e^{j})^{\top}\pm e^{j}(e^{i})^{\top}$. Then, the following statements hold:
  1. (i)
    If $f$ is spherically convex, then $b_{k}\leq 2(\delta_{ik}+\delta_{jk}-1)$ for all $k\in[n]$.
  2. (ii)
    If $b_{k}\leq-3\mp 1$, for all $k\in[n]$, then $f$ is spherically convex.
Proof.
  1. (i)
    Let $f$ be spherically convex. If $k\notin\{i,j\}$, then the statement follows by using inequality (1) with $u=(1/\sqrt{2})e^{i}\mp(1/\sqrt{2})e^{j}$ and $v=e^{k}$. If $k\in\{i,j\}$, then the inequality follows by using inequality (1) with $u=e^{\ell}$ and $v=e^{k}$, where $\ell\in\{i,j\}\setminus\{k\}$.
  2. (ii)
    Let $b_{k}\leq-3\mp 1$, for all $k\in[n]$. First, suppose that $A=e^{i}(e^{j})^{\top}+e^{j}(e^{i})^{\top}$. Let any $u\in\mathbb{S}^{n}$ and any $v\in\mathbb{S}^{n}\cap K$ with $\left\langle u,v\right\rangle=0$. We have
    $\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle=2u_{i}u_{j}-2v_{i}v_{j}\geq-(u_{i}^{2}+u_{j}^{2})-(v_{i}^{2}+v_{j}^{2})\geq-\sum_{k=1}^{n}u_{k}^{2}-\sum_{k=1}^{n}v_{k}^{2}$
    $=-2\sum_{k=1}^{n}v_{k}^{2}\geq-2\sum_{k=1}^{n}v_{k}=\frac{1}{2}\left(-4\sum_{k=1}^{n}v_{k}\right)\geq\frac{1}{2}\left(\sum_{k=1}^{n}b_{k}v_{k}\right)=\frac{1}{2}\left\langle b,v\right\rangle,$
    which implies inequality (1). Therefore, $f$ is spherically convex. Next, suppose that $A=-e^{i}(e^{j})^{\top}+e^{j}(e^{i})^{\top}$. Let any $u\in\mathbb{S}^{n}$ and any $v\in\mathbb{S}^{n}\cap K$ with $\left\langle u,v\right\rangle=0$. We have
    $\left\langle Au,u\right\rangle-\left\langle Av,v\right\rangle=2v_{i}v_{j}-2u_{i}u_{j}\geq-2u_{i}u_{j}\geq-(u_{i}^{2}+u_{j}^{2})\geq-\sum_{k=1}^{n}u_{k}^{2}$
    $=-\sum_{k=1}^{n}v_{k}^{2}\geq-\sum_{k=1}^{n}v_{k}=\frac{1}{2}\left(-2\sum_{k=1}^{n}v_{k}\right)\geq\frac{1}{2}\left(\sum_{k=1}^{n}b_{k}v_{k}\right)=\frac{1}{2}\left\langle b,v\right\rangle,$
    which implies inequality (1). Therefore, $f$ is spherically convex.
The following theorem and corollary provide additional large classes of spherically convex functions.
Theorem 7.
Let $K=\mathbb{R}^{n}_{+}$,
$b=\sum_{i=1}^{n}a_{ii}^{+}b^{i}_{+}+\sum_{i=1}^{n}a_{ii}^{-}b^{i}_{-}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{+}b_{+}^{ij}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{-}b_{-}^{ij},$
where $b^{i}_{+},b^{i}_{-},b^{ij}_{+},b^{ij}_{-}\in\mathbb{R}^{n}$ with ${b^{i}_{+}}_{k},{b^{i}_{-}}_{k}\leq-2\delta_{ik},$ ${b^{ij}_{-}}_{k}\leq-2$, and ${b^{ij}_{+}}_{k}\leq-4$, for all $i,j,k\in[n]$. Then, $f$ is spherically convex.
Proof.
Since
$A=\sum_{i=1}^{n}a_{ii}^{+}e_{i}e_{i}^{\top}+\sum_{i=1}^{n}a_{ii}^{-}\left(-e_{i}e_{i}^{\top}\right)+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{+}\left(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}\right)+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{-}\left(-e_{i}e_{j}^{\top}-e_{j}e_{i}^{\top}\right)$
and $f=f_{A,b,c}$, we have
$f=\sum_{i=1}^{n}a_{ii}^{+}f_{e_{i}e_{i}^{\top},b^{i}_{+},c^{i}_{+}}+\sum_{i=1}^{n}a_{ii}^{-}f_{-e_{i}e_{i}^{\top},b^{i}_{-},c^{i}_{-}}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{+}f_{e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top},b^{ij}_{+},c^{ij}_{+}}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{-}f_{-e_{i}e_{j}^{\top}-e_{j}e_{i}^{\top},b^{ij}_{-},c^{ij}_{-}},$
where $c^{i}$, $c^{ij}_{+}$ and $c^{ij}_{-}$ are arbitrary real constants with
$c=\sum_{i=1}^{n}a_{ii}c^{i}_{+}+\sum_{i=1}^{n}a_{ii}c^{i}_{-}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{+}c_{+}^{ij}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{-}c_{-}^{ij}.$
Hence, items (ii) of Lemmas 4-5 imply that $f$ is a linear combination of spherically convex functions with nonnegative coefficients. Therefore, $f$ is spherically convex. ∎
Corollary 4.
Suppose that $K=\mathbb{R}^{n}_{+}$ and $A\neq 0$. Then, the following statements hold:
  1. (i)
    Let
    $b=\sum_{i=1}^{n}|a_{ii}|b^{i}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}|a_{ij}|b^{ij},$
    where $b^{i},b^{ij}\in\mathbb{R}^{n}$ with $b^{i}_{k}\leq-2\delta_{ik}$ and $b^{ij}_{k}\leq-4$, for all $i,j,k\in[n]$. Then, $f$ is spherically convex.
  2. (ii)
    If
    $b_{k}\leq-2|a_{kk}|-4\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{+}-2\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{-},$
    for all $k\in[n]$, then $f$ is sperically convex.
  3. (iii)
    If
    $b_{k}\leq-2|a_{kk}|-4\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}|a_{ij}|,$
    for all $k\in[n]$, then $f$ is sperically convex.
Proof.
  1. (i)
    Since $|a_{ij}|=a_{ij}^{+}+a_{ij}^{-}$, for all $i,j\in[n]$, the result follows from Theorem 7 with $b^{i}_{+}=b^{i}_{-}=b^{i}$ and $b^{ij}_{+}=b^{ij}_{-}=b^{ij}$.
  2. (ii)
    Since $|a_{ii}|=a_{ii}^{+}+a_{ii}^{-}$, for all $i\in[n]$, the result follows from Theorem 7 with
    $b^{i}_{+k}=b^{i}_{-k}=-2\delta_{ik}\frac{b}{\displaystyle-2\sum_{i=1}^{n}|a_{ii}|-4\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{+}-2\delta{ik}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{-}},$
    $b^{ij}_{-}=-2\frac{b}{\displaystyle-2\sum_{i=1}^{n}|a_{ii}|-4\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{+}-2\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}a_{ij}^{-}}$
    and $b^{ij}_{+}=2b^{i}_{+}$, for all $i,j,k\in[n]$.
  3. (iii)
    It follows from $|a_{ij}|=a_{ij}^{+}+a_{ij}^{-}$, for all $i,j\in[n]$ and item ii.
5 Final remarks
In this paper we studied the geodesic convexity of non-homogeneous quadratic functions defined on a proper spherically convex set (i.e., a geodesically convex set on the sphere), which we simply called spherical convexity. We presented various necessary, sufficient and equivalent conditions for the spherical convexity of such functions. Besides being useful for positive and negative certificates regarding their spherical convexity, these conditions also yield intriguing consequences.
  1. 1.
    For any symmetric matrix that defines the homogeneous quadratic part of a non-homogeneous quadratic function, there exist infinitely many vectors that define the homogeneous linear term of the function, ensuring that the function is spherically convex, making the class of spherically convex non-homogeneous quadratic functions much larger than the class of (Euclidean) convex non-homogeneous quadratic functions.
  2. 2.
    The homogeneous linear term of the function has no infuence on the (Euclidean) convexity of the function, but it is crucial for its spherical convexity.
  3. 3.
    Even more interestingly, the previous consequence sharply contrasts with the fact that the class of spherically convex homogeneous functions on the spherically convex set formed by unit vectors of positive coordinates is formed by the constant functions, hence it is much smaller than its Euclidean counterpart.
We found specific explicit conditions for the spherical convexity of non-homogeneous quadratic functions correspondig to positive matrices, Z-matrices and diagonal matrices.
A challenge for the future is to find explicit conditions for more general non-homogeneous quadratic functions, defined on more general spherically convex sets and/or corresponding to more general matrices. It would be particularly interesting to study the spherical convexity of non-homogeneous quadratic functions on the spherically convex set defined by the intersection of the Lorentz cone with the sphere.